# FPT+: A Parameter and Memory Efficient Transfer Learning Method for High-resolution Medical Image Classification

### v1

@@ Abstract
The success of large-scale pre-trained models has established fine-tuning as a standard method for achieving significant improvements in downstream tasks. However, fine-tuning the entire parameter set of a pre-trained model is costly. Parameter-efficient transfer learning (PETL) has recently emerged as a cost-effective alternative for adapting pre-trained models to downstream tasks. Despite its advantages, the increasing model size and input resolution present challenges for PETL, as the training memory consumption is not reduced as effectively as the parameter usage. In this paper, we introduce Fine-grained Prompt Tuning plus (FPT+), a PETL method designed for high-resolution medical image classification, which significantly reduces memory consumption compared to other PETL methods. FPT+ performs transfer learning by training a lightweight side network and accessing pre-trained knowledge from a large pre-trained model (LPM) through fine-grained prompts and fusion modules. Specifically, we freeze the LPM and construct a learnable lightweight side network. The frozen LPM processes high-resolution images to extract fine-grained features, while the side network employs the corresponding down-sampled low-resolution images to minimize the memory usage. To enable the side network to leverage pre-trained knowledge, we propose fine-grained prompts and fusion modules, which collaborate to summarize information through the LPM’s intermediate activations. We evaluate FPT+ on eight medical image datasets of varying sizes, modalities, and complexities. Experimental results demonstrate that FPT+ outperforms other PETL methods, using only 1.03% of the learnable parameters and 3.18% of the memory required for fine-tuning an entire ViT-B model. Our code is available online.
keywords:
@@
Parameter-efficient transfer learning , Memory-efficient transfer learning , Large-scale pre-trained model , High-resolution medical image classification
journal: Arxiv
[…]
1 Introduction
@@
Utilizing the technique of transfer learning (Weiss et al., 2016), pre-trained models can be effectively adapted to specific downstream tasks by initializing task-specific models with weights from a pre-trained model, followed by training on task-specific datasets. The landscape has notably been reshaped by the remarkable achievements of large-scale pre-trained models (LPMs) across diverse domains (Radford et al., 2019; Devlin et al., 2019; Yang et al., 2019; Radford et al., 2021; Zhang et al., 2023b; Kirillov et al., 2023; Cheng et al., 2023). As model sizes grow rapidly (e.g. ViT-B, a variant of vision transformer (Dosovitskiy et al., 2021), has 86 million parameters), fine-tuning the entire parameter set of an LPM has become very costly.
@@
Figure 1: We utilize vision transformer (ViT-B) (Dosovitskiy et al., 2021), pre-trained on ImageNet-21K (Ridnik et al., 2021), as the encoder. The reported performance is the average of performance metrics across eight medical image datasets. The size of the dot denotes the memory usage during transfer learning. Our proposed method, FPT+, distinguishes itself from other PETL methods by achieving the best balance between performance and efficiency.
@@
In response, the concept of parameter-efficient transfer learning (PETL) (Ding et al., 2023) has emerged. PETL offers a strategic approach for transferring pre-trained models by selectively updating a small subset of the pre-trained parameters or introducing a modest number of additional parameters specific to the new task while freezing the majority of the pre-trained parameters.
@@
PETL initially gained substantial attention within the natural language processing (NLP) community before finding its application in computer vision (CV) where various PETL methods have shown effectiveness across diverse tasks and datasets. While these approaches have achieved notable success, they were primarily designed for natural images and the field of medical image analysis has not yet fully benefited from such advances (Dutt et al., 2023). The unique challenges of medical image analysis require specialized PETL methods due to the differences between natural and medical images (Shen et al., 2017; Zu et al., 2024; Kim et al., 2024).
@@
In natural images, the objects of interest typically occupy a significant portion of the image and have clear and noticeable attributes (e.g., shape and color). In contrast, the diagnostic cues within medical images are often tiny and spread throughout the entire visual field, making fine-grained information crucial for accurate medical image analysis (Huang et al., 2024b). Providing such fine-grained information often requires high-resolution input images (Huang et al., 2023). However, as shown in Fig. 2, this preference for high resolution comes at the cost of increased GPU memory consumption and higher training expenses. For instance, in a vision transformer (ViT) (Dosovitskiy et al., 2021) with a patch size of 16, a standard 224-resolution image is divided into 196 patches, while a 512-resolution image generates 1024 patches. Such high computational demand makes fine-tuning LPMs expensive and, in some cases, infeasible on devices with limited GPU memory.
@@
Furthermore, it is important to note that existing LPMs are primarily pre-trained on large natural image datasets like ImageNet (Deng et al., 2009). Attempting to bridge the knowledge gap between natural and medical images by merely updating a small subset of the pre-trained parameters or introducing additional parameters within a frozen LPM may not be sufficient to address the significant differences between these two domains (Zu et al., 2024).
@@
In this work, we propose a novel parameter and memory-efficient PETL method, namely Fine-grained Prompt Tuning plus (FPT+). FPT+ aims to enhance the efficiency of PETL specifically for high-resolution medical images. Existing PETL methods typically involve training a subset of parameters within the LPM of interest. However, updating these parameters requires gradients’ computation through the frozen layers during back-propagation, significantly increasing memory consumption. To address this issue, FPT+ employs a lightweight additive network inspired by the concept of a side network (Zhang et al., 2020; Sung et al., 2022). This design treats the entire pre-trained model as a frozen feature extractor, eliminating the need for back-propagation through the LPM. To leverage pre-trained knowledge from the LPM, FPT+ introduces the concept of fine-grained prompts and Fine-grained Fusion Modules (FFMs) as the bridging components. Fine-grained prompts consist of a small set of trainable embeddings prepended to the input sequence of the side network. Instead of engaging the entire input sequence for interaction with the fine-grained features from the pre-trained model, the fusion process involves only the fine-grained prompts through cross-attention mechanisms. These prompts subsequently convey fine-grained information to the side network during forward propagation, allowing the integration of pre-trained knowledge into the learnable side network at a low cost.
@@
Additionally, considering that long input sequences from high-resolution images result in substantial intermediate activations, which can be computationally expensive, FPT+ employs important tokens selection to mitigate this issue. Based on the prior knowledge that medical images of the same modality often share similar anatomical characteristics and have small objects of interest (e.g., lesions and vessels in fundus images), FPT+ selectively fuses features from the intermediate activations to reduce GPU memory consumption. Preloading techniques are also proposed to improve the training efficiency by pre-storing the intermediate features of the LPM.
@@
  • 1.
    We propose a novel PETL method, namely Fine-grained Prompt Tuning plus (FPT+), for high-resolution medical image classification. FPT+ maximizes the memory efficiency of transfer learning through asymmetric input and important tokens selection. The effectiveness and efficiency of adapting out-of-domain knowledge are achieved by our fine-grained prompts and fusion modules.
  • 2.
@@
  • 3.
    Extensive experiments are conducted on eight medical image datasets with various sizes, modalities and complexities, demonstrating the effectiveness of FPT+ by achieving the best balance between performance and parameter/memory efficiency.
  • 4.
@@
This paper extends our previous work FPT (Huang et al., 2024a). We systematically investigate the impact of several key components in FPT and refine the entire framework to make it simpler and more efficient. Specifically, our extension is presented in several aspects: (1) We reduce the number of layers in the side network and its corresponding fusion modules to significantly decrease memory consumption; (2) We redesign the fusion module by replacing in-projector with layer normalization and align the initial hidden dimension of the fine-grained prompts with the frozen LPM; (3) The new design improves the effectiveness of learning pre-trained knowledge, allowing smaller low-resolution inputs for the side network, thus further reducing memory consumption; (4) To verify the generalizability of FPT+, we validate the framework using four additional datasets; (5) We evaluate the scalability and feasibility of FPT+ across pre-trained models of different scales and different weights; (6) Extensive ablation studies are conducted to qualitatively and quantitatively analyze the impact of each component in FPT+ on performance, parameter usage and memory efficiency.
2 Related Work
2.1 Parameter-efficient Transfer Learning
@@
In recent years, the practice of fine-tuning expansive pre-trained models has shown encouraging potential across a range of tasks in the field of NLP. With the rapidly increasing sizes of pre-trained models, typically transformers (Vaswani et al., 2017; Brown et al., 2020), the concept of PETL has emerged as a strategic solution to mitigate the computational burden in adapting large-scale pre-trained models to downstream tasks.
@@
LoRA-based methods (Hu et al., 2022; Zhang et al., 2023a; YEH et al., 2024) freeze pre-trained weights and inject trainable rank decomposition matrices into each layer of a transformer to reduce the number of trainable parameters. Prompt tuning (Lester et al., 2021) appends the input with a trainable tensor, termed prompt, while freezing all other parameters. Adapter (Houlsby et al., 2019) introduces the idea of adding new learnable modules among the layers of the frozen pre-trained model to adapt to new downstream tasks. BitFit (Ben-Zaken et al., 2022) tunes only the bias term of a pre-trained model to effectively adapt the model to downstream tasks.
@@
Drawing inspiration from advancements in NLP, large-scale pre-trained models for CV are also receiving growing attention. ViTs are widely adopted as the backbone due to their modeling power and scalability to model and data size (Zhang et al., 2023a; Zhou et al., 2023; Ma et al., 2024; Cai et al., 2024). While most PETL methods were initially designed for transformers, recent works have demonstrated their effectiveness in ViTs as well. There also exist several PETL methods specifically designed for ViTs (Chen et al., 2022; Jia et al., 2022; Zhou et al., 2022b; Liu et al., 2023; Gong et al., 2023).
@@
Attention tuning (Touvron et al., 2022) fine-tunes only the attention layers to adapt ViTs to various downstream tasks. Vision prompt tuning (Jia et al., 2022) expands upon the concept of prompt tuning by introducing learnable prompts into each block of a frozen pre-trained ViT. Side tuning (Zhang et al., 2020) employs an additive side network, which combines features from a frozen pre-trained model using a simple summation operation facilitated by a trainable gate. Liu et al. (Liu et al., 2023) propose Explicit Visual Prompting to improve fine-tuning on segmentation tasks by taking the high-frequency components in the input as the prompts. CoOp (Zhou et al., 2022b) and CoCoOp (Zhou et al., 2022a) are proposed to specifically adapt CLIP-like vision-language pre-trained models to downstream image recognition tasks. Zhao et al. (Zhao et al., 2024) leverage task-specific information to select salient channels and then fine-tune only those channels to enhance parameter efficiency.
@@
Unlike other PETL methods, FPT (Huang et al., 2024a) and our proposed FPT+ aim to improve both parameter and memory efficiency during transfer learning, which is a specific and common requirement in the field of medical image analysis.
2.2 PETL for Medical Image Analysis
@@
While several PETL methods have emerged in the realm of CV, the application of PETL to ViTs for medical images remains relatively limited, despite the efficiency requirement inherent in the medical domain (Dutt et al., 2023). Raman et al. (Dutt et al., 2023) comprehensively evaluate the performance of different PETL methods for medical image analysis. Rodríguez et al. (Silva-Rodriguez et al., 2023) study adapters designed for dense prediction tasks to improve medical image segmentation. Dynamic Visual Prompt Tuning (He et al., 2023a) is designed to learn sample-specific and domain-specific features to enhance the fine-tuning performance in the medical domain. Zu et al. (Zu et al., 2024) propose Embedded Prompt Tuning by embedding prompt tokens into expanded channels for few-shot medical image classification tasks. Wu et al. (Wu et al., 2023) combine the SAM model (Kirillov et al., 2023) with adapters, which outperforms existing SOTA methods on several downstream medical image segmentation tasks.
@@
Our previous work, FPT (Huang et al., 2024a), to the best of our knowledge, is the first to enhance the efficiency of PETL in high-resolution input settings. We extend FPT by simplifying the entire framework to make it more efficient and more effective. Specifically, we halve the number of layers in the side network and the corresponding fusion modules, and redesign the fusion module’s in-projector. Unlike other PETL methods for medical image analysis, our proposed FPT+ framework mainly aims to address the challenge of unaffordable memory consumption in high-resolution medical image classification tasks.
@@
3 Method
In this work, we propose FPT+, a memory-efficient PETL method for high-resolution medical image classification. FPT+ aims to: 1) efficiently integrate fine-grained information from a pre-trained model into a new downstream model; and 2) effectively adapt knowledge from an out-of-domain pre-trained model. As shown in Fig. 3, a separate and lightweight side network is first employed to learn and predict at an affordable training cost in terms of both the number of trainable parameters and memory consumption. Then, fine-grained prompts and fusion modules are proposed to convey fine-grained pre-trained knowledge to the side network effectively and efficiently. Asymmetric input and important tokens selection are introduced to further reduce the computational burden. The motivation, benefits, and functioning of each component are delineated in subsequent sections.
3.1 Vision Transformer
In this section, we provide a brief overview of ViT (Dosovitskiy et al., 2021). A conventional ViT architecture consists of a linear projection layer followed by multiple transformer blocks. For a given image $x\in\mathbb{R}^{H\times W\times C}$ with spatial dimensions $(H,W)$ and channel dimension $C$, the initial step involves partitioning the image into non-overlapping 2D patches of size $P\times P$. Here, $P$ represents the patch size specific to the ViT under consideration. Subsequently, these 2D patches are flattened and transformed into a latent embedding space with hidden dimension $D$, facilitated by a linear projection layer. This process yields an input sequence $\bm{z}_{0}={z_{p}^{i}\in\mathbb{R}^{D}|i=1,...,N}$ with a length of $N=HW/P^{2}$. Each element in this input sequence is termed a token. The input is then processed by multiple transformer blocks, each consisting of a multi-headed self-attention layer (MSA) and a multi-layer perceptron block (MLP):

### v2

# Boosting Memory Efficiency in Transfer Learning for High-Resolution Medical Image Classification

@@ Abstract
The success of large-scale pre-trained models has established fine-tuning as a standard method for achieving significant improvements in downstream tasks. However, fine-tuning the entire parameter set of a pre-trained model is costly. Parameter-efficient transfer learning (PETL) has recently emerged as a cost-effective alternative for adapting pre-trained models to downstream tasks. Despite its advantages, the increasing model size and input resolution present challenges for PETL, as the training memory consumption is not reduced as effectively as the parameter usage. In this paper, we introduce Fine-grained Prompt Tuning plus (FPT+), a PETL method designed for high-resolution medical image classification, which significantly reduces the training memory consumption compared to other PETL methods. FPT+ performs transfer learning by training a lightweight side network and accessing pre-trained knowledge from a large pre-trained model (LPM) through fine-grained prompts and fusion modules. Specifically, we freeze the LPM of interest and construct a learnable lightweight side network. The frozen LPM processes high-resolution images to extract fine-grained features, while the side network employs corresponding down-sampled low-resolution images to minimize the memory usage. To enable the side network to leverage pre-trained knowledge, we propose fine-grained prompts and fusion modules, which collaborate to summarize information through the LPM’s intermediate activations. We evaluate FPT+ on eight medical image datasets of varying sizes, modalities, and complexities. Experimental results demonstrate that FPT+ outperforms other PETL methods, using only 1.03% of the learnable parameters and 3.18% of the memory required for fine-tuning an entire ViT-B model. Our code is available online.
Index Terms:
@@
Parameter-efficient transfer learning, Memory-effificent transfer learning, Large-scale pre-trained models, High-resolution medical image classification.
I Introduction
@@
Recently, large pre-trained models (LPMs) have made remarkable achievements across diverse domains [38, 13, 49, 37, 52, 29, 8]. Utilizing the technique of transfer learning [47], pre-trained models can be effectively adapted to specific downstream tasks by initializing task-specific models with weights from a pre-trained model of interest, followed by training on task-specific datasets. However, as model sizes grow rapidly (e.g. ViT-B, a variant of vision transformer (ViT) [15], has 86 million parameters), fine-tuning the entire parameter set of an LPM has become very costly, demanding substantial training resources such as GPU memory and significantly limiting its practical applicability.
@@
Figure 1: We utilize vision transformer (ViT-B) [15], pre-trained on ImageNet-21K [40], as the encoder. The reported performance is the average of performance metrics across eight medical image datasets. The size of the dot denotes the memory consumption during transfer learning. Our proposed method, FPT+, distinguishes itself from other PETL methods by achieving the best balance between performance and efficiency.
@@
In response, the concept of PETL [14] has emerged. PETL offers a strategic approach for transferring pre-trained models by selectively updating a small subset of the pre-trained parameters or introducing a modest number of additional parameters specific to the new task while freezing the majority of the pre-trained parameters.
@@
PETL initially gained substantial attention within the natural language processing (NLP) community before finding its application in computer vision (CV) where various PETL methods have shown effectiveness across diverse tasks and datasets. While these approaches have achieved notable success, they were primarily designed for natural images and the field of medical image analysis has not yet fully benefited from such advances [16]. The unique challenges of medical image analysis require specialized PETL methods due to the domain gap between natural and medical images [41, 57, 28].
@@
In natural images, the objects of interest typically occupy a significant portion of the image and exhibit clear and noticeable attributes (e.g., shape and color). In contrast, the diagnostic cues within medical images are often tiny and spread throughout the entire visual field, making fine-grained information crucial for accurate medical image analysis [25]. Providing such fine-grained information often requires high-resolution input images [24]. However, as illustrated in Fig. 2, this preference for high resolution comes at the cost of increased GPU memory consumption and higher training expenses. For example, a ViT with a patch size of 16 divides a 224 $\times$ 224 image into 196 patches, whereas a 512 $\times$ 512 image generates 1024 patches. This sharp rise in computational demand makes fine-tuning LPMs both expensive and, in some cases, infeasible on devices with limited GPU memory.
@@
Moreover, it is worth noting that existing LPMs are primarily pre-trained on large natural image datasets like ImageNet [12]. Attempting to bridge the knowledge gap between natural and medical images by merely updating a small subset of the pre-trained parameters or introducing additional parameters within a frozen LPM may not be sufficient to address the significant differences between these two domains [57].
@@
In this work, we propose a novel parameter and memory-efficient PETL method, namely Fine-grained Prompt Tuning plus (FPT+). FPT+ aims to enhance the efficiency of PETL specifically for high-resolution medical images. Existing PETL methods typically involve training a subset of parameters within the LPM of interest. However, updating these parameters still requires gradients’ computation through the frozen layers during back-propagation, significantly increasing memory consumption. To address this issue, FPT+ employs a lightweight additive network inspired by the concept of a side network [50, 44]. This design treats the entire pre-trained model as a frozen feature extractor, eliminating the need for back-propagation through the LPM. To leverage pre-trained knowledge from the LPM, FPT+ introduces the concept of fine-grained prompts and fine-grained fusion modules as the bridging components. Fine-grained prompts consist of a small set of trainable embeddings prepended to the input sequence of the side network. Instead of engaging the entire input sequence for interaction with the fine-grained features from the pre-trained model, the fusion process involves only the fine-grained prompts through cross-attention mechanisms. These prompts subsequently convey fine-grained information to the side network during forward propagation, allowing the integration of pre-trained knowledge into the learnable side network at a low cost.
@@
Additionally, considering that long input sequences from high-resolution images result in substantial intermediate activations, which can be computationally expensive, FPT+ employs important tokens selection to mitigate this issue. Based on the prior knowledge that medical images of the same modality often share similar anatomical characteristics and have small objects of interest (e.g., lesions and vessels in fundus images), FPT+ selectively fuses features from the intermediate activations to reduce GPU memory consumption. Preloading techniques are also proposed to improve the training efficiency by pre-storing the LPM’s intermediate features.
@@
  • We propose a novel PETL method, namely FPT+, for high-resolution medical image classification. FPT+ maximizes the memory efficiency of transfer learning through asymmetric input and important tokens selection. The effectiveness and efficiency of adapting out-of-domain knowledge are achieved by our fine-grained prompts and fusion modules.
@@
  • Extensive experiments are conducted on eight medical image datasets with various sizes, modalities and complexities, demonstrating the effectiveness of FPT+ by achieving the best balance between performance and parameter/memory efficiency. Additionally, resource consumption is comprehensively evaluated to establish the superior training efficiency of FPT+.
@@
This paper extends our previous work FPT [23]. We systematically investigate the impact of several key components in FPT and refine the entire framework to make it simpler and more efficient. Specifically, our extension is presented in several aspects: (1) We reduce the number of layers in the side network and its corresponding fusion modules to significantly decrease memory consumption; (2) We redesign the fusion module by replacing in-projector with layer normalization and align the initial hidden dimension of the fine-grained prompts with the frozen LPM; (3) All fusion modules share the same weights for the out-projector, further reducing the number of learnable parameters; (4) The new design improves the effectiveness of learning pre-trained knowledge, allowing smaller low-resolution inputs for the side network, thus further reducing memory consumption; (5) To verify the generalizability of FPT+, we validate the framework using four additional datasets; (6) We evaluate the scalability and feasibility of FPT+ across pre-trained models of different scales and different weights; (7) Additional efficiency metrics, including training time, throughput, and power usage, are evaluated to comprehensively assess FPT+’s resource consumption and training efficiency; (8) Extensive ablation studies are conducted to qualitatively and quantitatively analyze the impact of each component in FPT+ on performance, parameter usage and memory efficiency.
II Related Work
II-A Parameter-efficient Transfer Learning
@@
In recent years, the practice of fine-tuning expansive pre-trained models has shown encouraging potential across a range of tasks in the field of NLP. With the rapidly increasing sizes of pre-trained models, typically transformers [46, 3], the concept of PETL has emerged as a strategic solution to mitigate the computational burden in adapting large-scale pre-trained models to downstream tasks.
@@
LoRA-based methods [22, 51] freeze pre-trained weights and inject trainable rank decomposition matrices into each layer of a transformer to reduce the number of trainable parameters. Prompt tuning [31] appends the input with a trainable tensor, termed prompt, while freezing all other parameters. Adapter [21] introduces the idea of adding new learnable modules among layers of a frozen pre-trained model to adapt to new downstream tasks. BitFit [2] tunes only the bias term of a pre-trained model to effectively adapt the model to downstream tasks.
@@
Drawing inspiration from NLP advancements, large-scale pre-trained models for CV are also receiving growing attention. ViTs are widely adopted as the backbone due to their modeling power and scalability to model and data size [51, 56, 35, 4]. While most PETL methods were initially designed for transformers, recent works have demonstrated their effectiveness in ViTs as well.
@@
There are also several PETL methods specifically designed for ViT [7, 27, 55, 34, 17]. Attention tuning [45] fine-tunes only the attention layers to adapt ViTs to various downstream tasks. Vision prompt tuning [27] expands upon the concept of prompt tuning by introducing learnable prompts into each block of a frozen pre-trained ViT. Side tuning [50] employs an additive side network, which combines features from a frozen pre-trained model using a simple summation operation facilitated by a trainable gate. Liu et al. [34] propose Explicit Visual Prompting to improve fine-tuning on segmentation tasks by taking the high-frequency components in the input as the prompts. CoOp [55] and CoCoOp [54] are proposed to specifically adapt CLIP-like vision-language pre-trained models to downstream image recognition tasks. Zhao et al. [53] leverage task-specific information to select salient channels and then fine-tune only those channels to enhance parameter efficiency.
@@
Unlike other PETL methods, FPT [23] and our proposed FPT+ aim to improve both parameter and memory efficiency during transfer learning, which is a specific and common requirement in the field of medical image analysis.
II-B PETL for Medical Image Analysis
@@
While several PETL methods have emerged in the realm of CV, the application of PETL to ViTs for medical images remains relatively limited, despite the efficiency requirement inherent in the medical domain [16]. Raman et al. [16] comprehensively evaluate the performance of different PETL methods for medical image analysis. Rodríguez et al. [43] study adapters designed for dense prediction tasks to improve medical image segmentation. Dynamic Visual Prompt Tuning [19] is designed to learn sample-specific and domain-specific features to enhance the fine-tuning performance in the medical domain. Zu et al. [57] propose Embedded Prompt Tuning by embedding prompt tokens into expanded channels for few-shot medical image classification tasks. Wu et al. [48] combine the SAM model [29] with adapters, which outperforms existing SOTA methods on several downstream medical image segmentation tasks.
@@
To the best of our knowledge, our previous work, FPT [23], is the first PETL method specifically designed for high-resolution input settings, simultaneously enhancing both parameter and memory efficiency. We extend FPT by simplifying the entire framework to make it more efficient and more effective. Specifically, we halve the number of layers in the side network and the corresponding fusion modules, and redesign the fusion module’s in-projector. Unlike other PETL methods for medical image analysis, our proposed FPT+ framework aims to simultaneously address the challenge of unaffordable parameter as well as memory utilization when transferring LPMs to high-resolution medical image classification tasks.
@@
III Method
In this work, we propose FPT+, a parameter and memory efficient transfer learning method for high-resolution medical image classification. FPT+ can: 1) efficiently integrate fine-grained information from a pre-trained model into a new downstream model; and 2) effectively adapt knowledge from an out-of-domain pre-trained model. As shown in Fig. 3, a separate and lightweight side network is first employed to learn and infer at an affordable training cost in terms of both the number of trainable parameters and memory consumption. Then, fine-grained prompts and fusion modules are proposed to convey fine-grained pre-trained knowledge to the side network effectively and efficiently. Asymmetric input and important tokens selection are introduced to further reduce the computational burden. The motivation, benefits, and functioning of each component are delineated in subsequent sections.
III-A Vision Transformer
In this section, we provide a brief overview of ViT [15]. A conventional ViT architecture consists of a linear projection layer followed by multiple transformer blocks. For a given image $x\in\mathbb{R}^{H\times W\times C}$ with spatial dimensions $(H,W)$ and channel dimension $C$, the initial step involves partitioning the image into non-overlapping 2D patches of size $P\times P$. Here, $P$ represents the patch size specific to the ViT of interest. Subsequently, these 2D patches are flattened and transformed into a latent embedding space with hidden dimension $D$, facilitated by a linear projection layer. This process yields an input sequence $\bm{z}_{0}={z_{p}^{i}\in\mathbb{R}^{D}|i=1,...,N}$ with a length of $N=HW/P^{2}$. Each element in this input sequence is termed a token. The input is then processed by multiple transformer blocks, each consisting of a multi-headed self-attention (MSA) layer and a multi-layer perceptron (MLP) block: